N2 + 3H2= 2NH3

If a sample of 3 mole of hydrogen gas (H2) was reacted, how many miles of ammonia (NH3) gas would be produced ?
1
Expert's answer
2020-03-31T12:33:05-0400

Solution:

The equation of a chemical reaction:

N2 + 3H2= 2NH

According to the chemical equation: n(N2) = n(H2)/3 = n(NH3)/2.

Consequently,

n(H2)/3 = n(NH3)/2;

n(NH3) = 2 * n(H2) / 3;

n(NH3) = (2 * 3) / 3 = 2 (moles).

n(NH3) = 2 moles of ammonia.


Answer: 2 moles of ammonia (NH3) gas would be produced.